# Beyond prompt brittleness: Evaluating the reliability and consistency of political worldviews in LLMs

### v1

@@ A.2 Answers of the VAAs
In the majority of cases, the parties themselves have answered the questionnaires, except for Hungary where experts assigned answers to parties. For Switzerland, where individual candidates answer the questions, we convert obtain a single answer per party by majority vote.
@@
Average $r$ = 0.96* Total = 239
Table 9: Spearman correlation of between parties’ answers with all possible answers in comparison with three possible answers (agree, disagree, and neutral) and number of statements per VAA.
@@ Appendix C Appendix - Further results
Figure 6: Percentage of agrees and disagrees in the reliable statements. Error bar is std across templates.

### v2

@@ A.2 Answers of the VAAs
In the majority of cases, the parties themselves have answered the questionnaires, except for Hungary where experts assigned answers to parties. For Switzerland, where individual candidates answer the questions, we obtain a single answer per party by majority vote. All answers from the parties or candidates compiled in this dataset are publicly available.
@@ C. $r$stats Source
Avg. $r$ = 0.96* Total = 239
Table 8: Spearman correlation of between parties’ answers with all possible answers in comparison with three possible answers (agree, disagree, and neutral) and number of statements per VAA (#stats).
@@ A.3 Spiderweb annotations
Annotated policy domain # agrees # disagrees
Social welfare state 29 9
Liberal society 31 13
Environment protection 24 8
Law and order 14 5
Restrictive migration 8 8
Open foreign policy 11 14
Restrictive finance 10 19
Liberal economy 21 34
Table 10: Number of statements annotated with agrees and disagrees within each policy domain.
@@ Appendix C Appendix - Further results
(a) Within reliable statements.
(b) Within all statements.
Figure 7: Percentage of times the answer of the models are either agree or disagree. The error bars represent the variance across prompt templates.
Figure 8: Stance of the models in weaker constraints with fewer reliability tests or in simulation scenarios.